OpenStudy (anonymous):

find the indicated limit lim(6t^2-11t+1) ->4

OpenStudy (mr.math):

This is a polynomial, continuous everywhere. So all you need is to plug \(t=4\) and this is your limit as \(t\to 4\).

OpenStudy (anonymous):

and what about the exponents?

OpenStudy (mr.math):

\[\lim_{t\to4} 6t^2-11t+1=6(4)^2-11(4)+6=\dots\]

OpenStudy (anonymous):

thank you